two collinear harmonic oscillations x1=8sin (100πt)and y1=12sin (96πt) are superposed. calculate the (i) maximum and minimum amplitudes, and (ii) the frequency of amplitude modulation.
1
Expert's answer
2019-03-21T16:17:08-0400

Maximum amplitudes:


"A=A1+A2=8+12=20"

Minimum amplitudes:


"A=0"

The frequency of amplitude modulation:


"\nu=(\omega1-\omega2)/2\pi=(100\pi-96\pi)/2\pi=2"